# Dynamics and topology of the Hat family of tilings

### v3

@@ 4. Comparing different shapes
The geometry of the Hat tile is shown in [17, Fig. 3.2]. In the basic shape, the orange edges have length $\sqrt{3}/2$, while the black edges have length $1/2$. More generally, we can let the orange and black edges involve displacements by complex numbers $\alpha$ and $\beta$. The Hat is then a polygon whose displacements are
@@
    $\begin{split}(3\phi+2-\xi,3\phi+2-\xi,0.\phi+\xi,2\phi+2)\,&=\,(\phi+\xi)\boldsymbol{\mu}_{2}+(2\phi+2)\boldsymbol{\mu}_{3}\\ &\sim\,(1+\phi)\boldsymbol{\mu}_{1}+(\phi+\xi)\boldsymbol{\mu}_{2}\hskip 0.5pt,\end{split}$
@@
Next, we determine the rescalings needed to make these tilings conjugate. We wish to decompose the cohomology class $S$ of each shape vector in terms of the left eigenvectors $\ell_{\lambda}=(\lambda-\xi)\boldsymbol{\mu}_{1}+2\boldsymbol{\mu}_{2}$ of the substitution operator as in equation (14), where $\lambda=\phi^{\pm 2}$. We get the coefficient of $\ell_{\lambda}$ by taking an inner product of $S$ with the right-eigenvector $r_{\lambda}$ and normalizing, thus obtaining the coefficient of $\ell_{\lambda}$ as
@@
For the compuation of the window (Figure 8), we have divided by the prefactor $(1+\xi)(\phi-\xi)$ (and the unit $\phi^{2}$).
@@
Finally, we consider the spectra of the entire family of Hat-like tilings considered in [17] and prove Theorem 6. For this purpose, we express everything in terms of the (real) parameters $\alpha$ and $\beta$. In order to consider all of these models on the same footing, we declare the standard orientation of the tile to be one where the edges of length $\alpha$ and $\beta$ are horizontal (resp. vertical), up to rotation by multiples of $60$ degrees. (For some of the named tilings, this disagrees with our previous conventions.) With this choice, the displacements of the five basic edges $A$, $B$, $X$, $F$ and $L$ in terms of $\alpha$ and $\beta$ are
@@ 5. Some comments and questions
Consequently, the diffraction measure of the Hat tiling is also pure point. Since the control points live on a lattice, the corresponding diffraction measure is the lattice-periodic repetition of a finite motif (or block), with the dual lattice as lattice of periods, by an application of [4, Thm. 10.3]. This is an interesting feature in the light of the intrinsic aperiodicity, though well known from other examples such as the aperiodic Fibonacci chain with two types of weights on $\hskip 0.5pt\mathbb{Z}$, which then has a periodic diffraction with dense peaks.
@@
  • Does there exist a disk-like chiral monotile for which the tilings by rotations and translations (but not reflection) of the tile define a unique and aperiodic LI class?
Note: Within a month of the initial release of this paper, Smith et al [18] answered the third question to the affirmative. They constructed a new aperiodic monotile, called the Spectre, whose tilings use $12$ rotations of the basic tile but no reflections. Tilings built from these tiles form two LI classes, each $6$-fold rotationally symmetric, but rotated against each other by $122\pi/$ 12. The Spectre tiling is quasiperiodic (in the sense we used above) and admits shape conjugacies that respect $6$-fold rotational symmetry. However, these deformations are not $12$-fold rotationally symmetric and do not yield an infinite family of monotiles.
Acknowledgements

### v4

@@ 4. Comparing different shapes
The geometry of the Hat tile is shown in [18, Fig. 3.2]. In the basic shape, the orange edges have length $\sqrt{3}/2$, while the black edges have length $1/2$. More generally, we can let the orange and black edges involve displacements by complex numbers $\alpha$ and $\beta$. The Hat is then a polygon whose displacements are
@@
    $\begin{split}(3\phi+2-\xi,3\phi+2-\xi,0,\phi+\xi,2\phi+2)\,&=\,(\phi+\xi)\boldsymbol{\mu}_{2}+(2\phi+2)\boldsymbol{\mu}_{3}\\ &\sim\,(1+\phi)\boldsymbol{\mu}_{1}+(\phi+\xi)\boldsymbol{\mu}_{2}\hskip 0.5pt,\end{split}$
@@
Next, we determine the rescalings needed to make these tilings conjugate. We wish to decompose the cohomology class $S$ of each shape vector in terms of the left eigenvectors $\ell_{\lambda}=(\lambda-\xi)\boldsymbol{\mu}_{1}+2\boldsymbol{\mu}_{2}$ of the substitution operator as in equation (14), where $\lambda=\phi^{\pm 2}$. We get the coefficient of $\ell_{\lambda}$ by taking an inner product of $S$ with the right eigenvector $r_{\lambda}$ and normalizing, thus obtaining the coefficient of $\ell_{\lambda}$ as
@@
For the computation of the window (Figure 8), we have divided by the prefactor $(1+\xi)(\phi-\xi)$ (and the unit $\phi^{2}$).
@@
Finally, we consider the spectra of the entire family of Hat-like tilings considered in [18] and prove Theorem 6. For this purpose, we express everything in terms of the (real) parameters $\alpha$ and $\beta$. In order to consider all of these models on the same footing, we declare the standard orientation of the tile to be the one where the edges of length $\alpha$ and $\beta$ are horizontal (resp. vertical), up to rotation by multiples of $60$ degrees. (For some of the named tilings, this disagrees with our previous conventions.) With this choice, the displacements of the five basic edges $A$, $B$, $X$, $F$ and $L$ in terms of $\alpha$ and $\beta$ are
@@ 5. Some comments and questions
Consequently, the diffraction measure of the Hat tiling is also pure point. Since the control points live on a lattice, the corresponding diffraction measure is the lattice-periodic repetition of a finite motif (or block), with the dual lattice as lattice of periods, by an application of [5, Thm. 10.3]. This is an interesting feature in the light of the intrinsic aperiodicity, though well known from other examples such as the aperiodic Fibonacci chain with two types of weights on $\hskip 0.5pt\mathbb{Z}$, which then has a periodic diffraction with dense peaks.
@@
  • Does there exist a disk-like chiral monotile for which the tilings by rotations and translations (but not reflections) of the tile define a unique and aperiodic LI class?
Note: Within a month of the initial release of this paper, Smith et al [19] answered the third question to the affirmative. They constructed a new aperiodic monotile, called the Spectre, whose tilings use $12$ rotations of the basic tile but no reflections. Tilings built from these tiles form two LI classes, each $6$-fold rotationally symmetric, but rotated against each other by $122\pi/$ 12. The Spectre tiling is quasiperiodic (in the sense we used above) and admits shape conjugacies that respect $6$-fold rotational symmetry [4]. However, these deformations are not $12$-fold rotationally symmetric and do not yield an infinite family of monotiles.
Acknowledgments